OpenStudy (theviper):

\(\Large{Tutorial:-}\) How to find the quadratic & cubic polynomials if there zeroes are given .....

OpenStudy (theviper):

QUADRATIC polynomial :- If zeroes are α & β then polynomial will be find by this :- \(x^2−x(α+β)+αβ\) CUBIC polynomial :- If zeroes are α,β & γ then polynomial will be find by this :- \(x^3−x^2(α+β+γ)+x(αβ+βγ+γα)−αβγ\)

OpenStudy (theviper):

Example1:- Find a quadratic polynomial, the sum & product of whose zeroes are -5 & 6 respectively. Solution:- Let \(\alpha\ \&\ \beta\) are the zeroes of the required polynomial \(f(x)\). Then, \(\alpha+\beta\)=-5 and \(\alpha \beta\)=6. \(\dot{\ddot{}}\)\(f(x)=x^2-x(\alpha+\beta )+\alpha \beta\) \(\implies f(x)=x^2-(-5x)+6\) \(\implies f(x)=x^2+5x+6.\) Hence, the required polynomial is \(f(x)=x^2+5x+6\)

OpenStudy (theviper):

Example2 :- Find a cubic polynomial whose zeroes are \(\alpha,\beta \) & \(\gamma\) such that \((\alpha+\beta+\gamma)\)=6, \((\alpha\beta+\beta\gamma+\gamma\alpha)\)=-1 & \((\alpha\beta\gamma)\)=-30. Solution :- We know that a cubic polynomial whose zeroes are \(\alpha\beta\) & \(\gamma\) is given by \(p(x)=x^3-x^2(\alpha+\beta+\gamma)+x(\alpha\beta+\beta\gamma+\gamma\alpha)\)-\(\alpha\beta\gamma\) \(\implies p(x)=x^3-6x^2-x-(-30)\) \(\implies p(x)=x^3-6x^2-x+30\) Hence, the required polynomial is \((x^3-6x^2-x+30)\).

OpenStudy (theviper):

@mathslover @vishweshshrimali5 @UnkleRhaukus I now gave examples as u siad :)

OpenStudy (compassionate):

What about a quadratic molynomial structure?

OpenStudy (theviper):

hmm... I m too small for that :D

OpenStudy (theviper):

LOL

OpenStudy (vishweshshrimali5):

One suggestion ........

OpenStudy (theviper):

Fast

OpenStudy (vishweshshrimali5):

Please increase the size of the latex

OpenStudy (theviper):

Same was given by @sasogeek :D

OpenStudy (vishweshshrimali5):

:D

OpenStudy (theviper):

LOL it is for help :P

OpenStudy (anonymous):

\[The \ Viper \\ \huge\ GOOD \ WORK\]

OpenStudy (theviper):

\[\Huge{\color{blue}{\cal{Thanx}}\color{green}{\ddot{\smile}}}\]@Rohangrr